Question 1 (8 marks)

An appliance that cost a retailer $1,800 less trade discounts of 37.5% and 18%, carries a price tag with a regular selling price at a markup of 120% of cost. For quick sale, the item was marked down 40%.

a.  What is the net price of the appliance? ___________

b.  What is the regular selling price? __________

c.  What is the sale price? __________

d.  What rate of markup based on cost was realized? ___________ 

Need it as soon as possible

1
Expert's answer
2021-02-17T03:09:39-0500

a)

"\frac{1800}{1-(0.375+0.18)}=4045~\text{}"$,

b)

"1800\cdot(1+1.20)=3960~" $,

c)

"3960\cdot(1-0.40)=2376~" $,

d)

"1-(1+1.20)\cdot(1-0.40)=0.32=32~"%.